# VRSD: Rethinking Similarity and Diversity for Retrieval in Large Language Models

### v1

@@ 4.2 Experiments Results
VRSD 0.371 0.233 0.789 0.534 0.213 0.198
MMR($\lambda$=0) 0.355 0.216 0.767 0.508 0.206 0.198
MMR($\lambda$=0.5) 0.364 0.218 0.772 0.507 0.202 0.188
MMR($\lambda$=1) 0.347 0.222 0.780 0.510 0.188 0.186
Table.1 presents the results of retrieval quality, indicating that the win rate of VRSD, compared to MMR across various datasets and conditions, consistently exceeds 90%. This outcome suggests that VRSD not only retrieves examples more relevant to the original query but also better satisfies the diversity requirements. Remarkably, VRSD maintains a minimum advantage of 0.05 over MMR in the worst-case scenario concerning the maximum difference. Additionally, the mean value of cosine similarity between the sum vector and the query vector is significantly higher for VRSD than for MMR. VRSD generally demonstrates superior retrieval efficacy from diverse perspectives. Table 2 shows the answer quality results, with VRSD achieving the highest scores across all metrics, thus indicating that the examples retrieved by VRSD enhance the LLM’s understanding of the query and facilitate the generation of the desired answers. The superior performance of the closed-source LLM over the open-source LLM can be attributed to the former’s enhanced understanding and reasoning capabilities. Overall, VRSD exhibits superior performance to MMR in both example retrieval and query answering, fulfilling the requirements of both similarity and diversity without the need for parameter adjustments, thereby highlighting the pronounced advantages of VRSD.
It is worth noting that, compared to other datasets, the mean values of $\cos(d_{\text{VRSD}},q)$ and $\cos(d_{\text{MMR}},q)$ in the Puzzle dataset are relatively low. Additionally, the difference in results between the open-source and closed-source models is less pronounced, and VRSD achieves a 100% win rate, which may be attributed to the small size of the dataset. Nevertheless, given that lateral-puzzle questions require LLMs to comprehend the query and derive insights from various angles in the retrieved examples, the Puzzle dataset remains a valuable tool for assessing our algorithm.
5 Conclusion
@@
In this work, given the complexity of parameter adjustment in MMR, we aim to enhance how LLMs retrieve similar and diverse examples by introducing a novel approach that characterizes both constraints through the relationship between the sum vector and the query vector. This method requires individual vectors within the sum to align divergently with the query vector, thereby satisfying the diversity constraint. We demonstrate that this problem is NP-complete, and we propose the VRSD algorithm, which not only outperforms MMR but also improves efficiency across various metrics, thereby enabling the retrieval of higher-quality examples. Our work underscores the inherent challenges of simultaneously pursuing similarity and diversity in vector retrieval and establishes a solid theoretical foundation for further research. The proposed combinatorial optimization problem holds independent interest from both theoretical and practical standpoints, indicating that further exploration and refinement of the heuristic algorithm would constitute a valuable avenue for future research. In terms of advancing the heuristic algorithm, we aim to explore two specific aspects: (1) the development of heuristic algorithms with lower time complexity that can retrieve high-quality examples more efficiently, and (2) the creation of adaptable heuristic algorithms that remain robust regardless of dataset size or problem type. We believe that as the retrieval algorithm—emphasizing both similarity and diversity—continues to improve, the scope of tasks that LLM-based agents can manage will expand, yielding increasingly satisfactory results.

### v2

@@ 4.2.1 Retrieval performance
VRSD 0.371 $\pm$ 0.008 0.233 $\pm$ 0.006 0.789 $\pm$ 0.004 0.534 $\pm$ 0.008 0.213 $\pm$ 0.007 0.198 $\pm$ 0.018
MMR($\lambda$=0) 0.355 $\pm$ 0.010 0.216 $\pm$ 0.004 0.767 $\pm$ 0.012 0.508 $\pm$ 0.014 0.206 $\pm$ 0.002 0.198 $\pm$ 0.015
MMR($\lambda$=0.5) 0.364 $\pm$ 0.011 0.218 $\pm$ 0.008 0.772 $\pm$ 0.006 0.507 $\pm$ 0.014 0.202 $\pm$ 0.002 0.188 $\pm$ 0.012
MMR($\lambda$=1) 0.347 $\pm$ 0.013 0.222 $\pm$ 0.005 0.780 $\pm$ 0.006 0.510 $\pm$ 0.014 0.188 $\pm$ 0.002 0.186 $\pm$ 0.011
Table.1 presents the overall result of MMR and VRSD on ARC-DA, OpenBookQA and Puzzle. From the results, we have the following observations and conclusions:
We observe that the win rate of VRSD consistently exceeds 90% compared to MMR across various datasets and conditions. This suggests that, without requiring additional parameters, VRSD retrieves examples that are more pertinent to the original query and more effectively satisfies diversity requirements. Notably, VRSD maintains an advantage over MMR in all scenarios concerning the Max-diff, underscoring VRSD’s superior capacity to leverage vector information for more diverse and relevant retrieval. Furthermore, MMR consistently underperforms across all datasets in terms of the Mean, which measures the average cosine similarity between the sum vector and the query vector. This finding indicates that the overall effectiveness of MMR retrieval is generally inferior to that of VRSD, and this discrepancy is not incidental. Additionally, the significant difference in Mean between VRSD and MMR suggests that the two methods capture different vectors when retrieving passages, further highlighting the distinction in their retrieval mechanisms.
4.2.2 Downstream task execution Performance
To verify that the improvements of VRSD in retrieval, with respect to both similarity and diversity, effectively enhance downstream tasks, we conducted a validation study using two models: the open-source LLM, Open-Mistral-7b, and the closed-source LLM, Gpt-3.5-Turbo. In this study, we reconstructed prompts by concatenating the original instances corresponding to the retrieved vectors with the initial query, and then input these prompts into the LLMs. The responses generated by the LLMs, based on different retrieval algorithms, were compared with standard answers using the ROUGE-L (Lin, 2004) metric for the ARC-DA and Puzzle datasets, and Accuracy (Schütze et al., 2008) for the OpenBookQA dataset, respectively. This evaluation allowed us to assess the efficacy of the retrieved examples in facilitating accurate responses. Furthermore, given that VRSD consistently outperformed MMR in retrieval performance across all models and datasets in previous results, we selected the model all-mpnet-base-v2 for retrieval prior to executing the downstream tasks.
The results in Table.2 shows that VRSD achieves the highest scores across all metrics for answer quality across all datasets. This suggests that the examples retrieved by VRSD better enhance the LLM’s understanding of the query and facilitate the generation of more accurate answers. Additionally, GPT-3.5-turbo outperforms open-mistral-7b, demonstrating that models with superior instruction-following capabilities benefit more from diverse and relevant examples. Moreover, as shown in Table.1, VRSD produces superior retrieval results, and these examples, when used by LLMs, lead to improved task performance. This not only confirms that the sum vector more effectively captures both diversity and similarity, but also that the retrieved examples substantially enhance the reliability of the LLM’s answers.
Overall, VRSD demonstrates superior performance compared to MMR in both retrieval effectiveness and task execution, meeting the demands of both similarity and diversity without requiring parameter adjustments. This underscores the distinct advantages of VRSD. Notably, in the Puzzle dataset, the mean values of $\cos(d_{\text{VRSD}},q)$ and $\cos(d_{\text{MMR}},q)$ are comparatively low and the discrepancy between open-source and closed-source models is less significant, with VRSD achieving a 100% win rate, likely due to the small dataset size. Nonetheless, since lateral puzzle questions necessitate that LLMs grasp the query and generate insights from multiple perspectives in the retrieved examples, the Puzzle dataset remains an important benchmark for evaluating our algorithm.
4.3 Further Analysis
Next, we perform a comprehensive analysis of BM25 and Cosine Similarity across all datasets to evaluate their performance in terms of diversity and similarity in retrieval. As BM25 and Cosine Similarity are classical distance measurement methods, it is crucial to assess their performance in relation to our sum vector; this enables us to understand their behavior within the context of the sum vector and justifies their inclusion in our comparison. While both BM25 and Cosine Similarity are rank-based retrieval approaches, VRSD distinguishes itself by simultaneously capturing both diversity and similarity, retrieving all relevant examples in one step and considering the potential connections between embedded vectors. For consistency, we select the same number of top-ranked examples for evaluation, similar to previous settings. As in prior experiments, we utilize the all-mpnet-base-v2 model for retrieval before executing downstream tasks.
Table.3 illustrates a notable performance decline when employing BM25 or direct cosine similarity for retrieval in terms of diversity and similarity, thereby confirming the superiority of VRSD. Furthermore, as reflected in the Mean metric, the retrieval performance follows the order: VRSD > Cosine Similarity > BM25. Similarly, the responses from both LLMs demonstrate the same downward trend. This finding suggests that VRSD more effectively balances diversity and relevance in the retrieved examples, ultimately enhancing the quality of LLM-generated responses. These insights from downstream tasks provide further evidence supporting our earlier assertion that the sum vector implicitly captures the diversity of the set while also accounting for similarity.
Table 3: Performance comparision of different algorithms on datasets. For each dataset, The first three indicators assess retrieval performance, while the latter two evaluate task execution performance. Win denotes VRSD win.rate.
Algorithm ARC-DA OpenBookQA Puzzle
Win Max-diff Mean GPT-3.5-turbo open-mistral-7b Win Max-diff Mean GPT-3.5-turbo open-mistral-7b Win Max-diff Mean GPT-3.5-turbo open-mistral-7b
VRSD - - $0.740^{\uparrow\uparrow}$ $0.371^{\uparrow\uparrow}$ $0.233^{\uparrow\uparrow}$ $0.833^{\uparrow\uparrow}$ $0.789^{\uparrow\uparrow}$ $0.534^{\uparrow\uparrow}$ $0.592^{\uparrow\uparrow}$ $0.213^{\uparrow\uparrow}$ $0.198^{\uparrow\uparrow}$
CS 95.3% 0.158 $0.710^{\uparrow}$ $0.361^{\uparrow}$ $0.222^{\uparrow}$ 96.8% 0.102 $0.812^{\uparrow}$ $0.780^{\uparrow}$ $0.510^{\uparrow}$ 100% 0.132 $0.577^{\uparrow}$ $0.201^{\uparrow}$ $0.187^{\uparrow}$
BM25 98.7& 0.684 0.550 0.347 0.196 99.9% 0.650 0.544 0.753 0.503 100% 0.262 0.459 0.188 0.185
5 Related Work
Retrieval methods in Large Language Models (LLMs) have gained traction, particularly due to their pivotal role in open-domain question answering, evidenced by seminal contributions in the field (Chen et al., 2017; Guu et al., 2020; Khattab & Zaharia, 2020; Qu et al., 2021; Izacard & Grave, 2021). The introduction of Retrieval Augmented Generation (RAG) further underscored the significance of these methods across knowledge-intensive tasks (Lewis et al., 2020), notably enhancing generation capabilities for open-domain queries (Mao et al., 2021). This spurred the adoption of techniques such as K-Nearest-Neighbor (KNN) in diverse applications, ranging from the customization of multilingual models in machine translation (Khandelwal et al., 2020) to improving the prediction of rare patterns in LLMs (Khandelwal et al., 2019; Alon et al., 2022). Continued advancements in retrieval techniques have focused on identifying highly informative examples to augment in-context learning, thereby enabling LLM-based systems to achieve significant performance improvements with minimal examples (Brown et al., 2020). Early on, traditional sparse retrieval methods like BM25 (Robertson et al., 2009)—an extension of TF-IDF—were utilized to refine in-context learning (Liu et al., 2022). Subsequently, the integration of LLMs’ intrinsic capabilities (Shin et al., 2021) and Sentence-BERT (SBERT) (Reimers & Gurevych, 2019) facilitated the retrieval of highly pertinent examples for prompt integration. The advent of dense retrievers signified a methodological enhancement in retrieval from a machine learning perspective, and the incorporation of feedback signals with contrastive learning has yielded more effective retrieval systems (Rubin et al., 2022; Wang et al., 2023). Recent innovations like UPRISE (Cheng et al., 2023) and PRAC (Nie et al., 2023) have further optimized the performance of in-context learning by retrieving demonstrations directly from training data. Despite these advances, most retrieval methods still treat each candidate independently, which can lead to suboptimal outcomes due to the interaction effects among in-context examples, resulting in a lack of diversity. Besides, given the expanding applications of LLMs, diversity becomes increasingly crucial, incorporating a diverse range of examples enriches LLMs’ learning processes, facilitating more innovative and robust responses, especially for complex open-ended questions. Most existing methods in this area rely on MMR; however, the necessity for continuous adjustment to determine the optimal value of $\lambda$ is inefficient in practical applications (Carbonell & Goldstein, 1998; Deselaers et al., 2009; Ye et al., 2023). In this work, we aim to retrieve examples from the perspective of combinatorial optimization, supported by strict theoretical analysis. Furthermore, the proposed sum vector has been demonstrated to effectively capture both diversity and similarity, contributing to the successful execution of downstream tasks.
6 Conclusions
@@
In this work, considering the complexity of parameter adjustment in MMR, we aim to improve how LLMs retrieve similar yet diverse examples by introducing a novel approach that jointly models both constraints through the relationship between the sum vector and the query vector. This method ensures that individual vectors within the sum diverge from the query vector, thereby satisfying the diversity constraint. We demonstrate that this problem is NP-complete and propose the VRSD algorithm, which not only surpasses MMR in retrieval performance but also enhances the execution of downstream tasks. Additionally, we introduce the sum vector, which is proven to effectively capture both diversity and similarity simultaneously, and serves as a metric to measure the balance between these two aspects, addressing an existing gap. Our work highlights the inherent challenges of achieving both similarity and diversity in vector retrieval, establishing a solid theoretical foundation for future research. The proposed combinatorial optimization problem holds independent value from both theoretical and practical perspectives, suggesting that further exploration and refinement of the heuristic algorithm would be a fruitful direction for future inquiry.